### v1

@@ 4.1 Auction Revenue and Market Efficiency
In this section, we examine the comparative statistics of equilibrium quantities that are relevant to sequencer design, starting with auction revenue.
@@
Conditional on a realization of the mixed strategy, the auctioneer takes the base gas and priority fee bid of the winner (if one exists), and $r_{1}$ times the base gas fee plus $r_{2}$ times the priority fee bid by the remaining participating arbitrageurs. Note that the revenue can be split into two terms: one capturing the revenue from the base gas fee, and the other capturing revenue from the priority gas fee bids. The expected revenue over all possible realizations along with its decomposition into base and priority components is characterized by the following theorem.
@@
If at least one of $r_{1}$ and $r_{2}$ is nonzero, under the symmetric mixed-strategy Nash equilibrium:
@@
Revenue is deeply connected with market efficiency. Recall that the probability that value from the arbitrage opportunity is extracted is $1-(p^{*})^{N}$. When this quantity is high, on-chain arbitrages are frequent, thereby improving market efficiency. On the other hand, when this quantity is low, on-chain arbitrages are less likely to occur, leaving arbitrage opportunities unexploited and generating zero value for the auction. It is straightforward to see that $r_{1}$ has the same directional impact on market efficiency (when measured by the probability that a value extraction occurs) as expected auction revenue described in Theorem 4.1, so more revert protection (lower $r_{1}$) implies a higher market efficiency.
@@ 4.2 Blockspace and Mempool Usage
Under the symmetric mixed-strategy Nash equilibrium:
@@ 4.3 Discussion: Full Revert Protection
4.3.1 Acknowledgements and Disclosures.
Ciamac C. Moallemi was supported by the Briger Family Digital Finance Lab at Columbia Business School and is an advisor to fintech companies.

### v2

@@ 4 Implications for Sequencer Design
In this section, we examine the comparative statistics of equilibrium quantities that are relevant to sequencer design.
@@ 4.1 Auction Revenue and Market Efficiency
Given a realization of the mixed strategy, the auctioneer takes the base fee and priority fee bid of the winner (if one exists), and $r_{1}$ times the base fee plus $r_{2}$ times the priority fee bids of the remaining participating arbitrageurs. The revenue can be split into two terms: one capturing the revenue from the base fee, and the other capturing revenue from priority fee bids. The expected revenue over all possible realizations along with its decomposition into base and priority components is characterized by the following theorem.
@@
If at least one of $r_{1}$ and $r_{2}$ is nonzero, then under the symmetric mixed-strategy Nash equilibrium:
@@
Revenue is deeply connected with market efficiency in automated market makers. Recall that the probability that value from the arbitrage opportunity is extracted is $1-(p^{*})^{N}$. When this quantity is high, on-chain arbitrages are frequent, thereby improving market efficiency. On the other hand, when this quantity is low, on-chain arbitrages are less likely to occur, leaving arbitrage opportunities unexploited and generating zero value for the auction. It is straightforward to see that $r_{1}$ has the same directional impact on market efficiency (when measured by the probability that a value extraction occurs) as expected auction revenue described in Theorem 4.1, so more revert protection (lower $r_{1}$) implies a higher market efficiency.
@@ 4.2 Blockspace and Mempool Usage
If at least one of $r_{1}$ and $r_{2}$ is nonzero, then under the symmetric mixed-strategy Nash equilibrium:
@@ 4.3 Discussion: Full Revert Protection
5 Extensions
So far, we assumed that revert protection can be implemented with negligible cost. This may not hold in practice, with block builders bearing non-negligible costs for implementing revert protection. We now take these costs into consideration and analyze two schemes in which costs are (i) internalized by the sequencer and (ii) passed on to searchers/arbitrageurs. In both scenarios, we now suppose that the cost of implementing RP per transaction submitted to the mempool is given by $c>0$. To avoid trivialities, we make the following assumption:
Assumption 5.1
Assume that the cost per transaction does not exceed the value less base fee, i.e., $c<V-g$.
Scheme 1: Sequencer Internalizes Costs.
Here, the sequencer incurs costs themselves. While the equilibrium for the searchers does not change, the sequencer’s expected revenue under internalization, denoted $\mathbb{E}[\textsf{Profit}]$, is
$\mathbb{E}[\textsf{Profit}]$ $=\mathbb{E}[\textsf{Revenue}]-c\cdot\mathbb{E}[\mathsf{SubmittedTXs}]=(1-(p^{*})^{N})V-c(1-p^{*})N.$
Since $p^{*}$ depends on $r_{1}$, the revert penalty rate on the base fee, we can optimize over $r_{1}$ to find the profit-optimal rate.
Theorem 5.1
If at least one of $r_{1}$ and $r_{2}$ is nonzero, then under the symmetric mixed-strategy Nash equilibrium, the profit-optimal $r_{1}$ is
$r_{1}^{*}=\frac{c(V-g)}{(V-c)g}\cdot\mathbbm{1}\{c\leq g\}+\mathbbm{1}\{c>g\}$
Unlike the baseline model, where full revert protection maximized revenue, a nonzero revert penalty rate now maximizes profits when sequencers internalize the costs of implementing RP. This result highlights a tradeoff between offering full RP to searchers and optimizing profit while covering the costs of full RP.
Scheme 2: Costs Passed Onto Searchers.
Here, searchers pay the builders an amount $c$ upfront for the cost of simulating their transactions. The cost $c$ can also represent a “cost of searching” incurred by searching for MEV opportunities. The searcher’s payoff is
$u_{i}(b_{i}|\beta_{-i})=\begin{dcases}(V-g-b_{i})\mathbb{P}(w=i|b_{i},\beta_{-i})&\\ \qquad-(r_{1}g+r_{2}b_{i})\!\left(1-\mathbb{P}(w=i|b_{i},\beta_{-i})\right)-c&\text{if $b_{i}\geq 0$,}\\ 0&\text{if $b_{i}=\varnothing$.}\end{dcases}$
We resolve for the equilibrium and implications for sequencer design under the modified payoffs.
Theorem 5.2
When searchers pay for the cost of revert protection
  • If $r_{1}=r_{2}=0$, then a pure strategy profile $b$ where bids are ordered such that $b_{1}\leq b_{2}\leq\dots\leq b_{N}$ (with abstentions represented by bids less than zero) is a Nash equilibrium if and only if $b_{N-1}=b_{N}=V-g-c$.
  • If at least one of $r_{1}$ and $r_{2}$ is nonzero, then a pure-strategy Nash equilibrium does not exist. The unique symmetric mixed-strategy equilibrium is given by
    $p_{i}^{*}=p^{*}\triangleq\left(\frac{r_{1}g+c}{V-g+r_{1}g}\right)^{\frac{1}{N-1}},$
    $F_{i}^{*}=F^{*}(b)\triangleq\frac{1}{1-p^{*}}\left(\left(\frac{r_{1}g+r_{2}b+c}{V-g-b+r_{1}g+r_{2}b}\right)^{\frac{1}{N-1}}-p^{*}\right),$
    for bids $b\in[0,V-g]$ and agents $i\in[N]$. The expected payoff of every agent in equilibrium is zero.
  • Expected revenue decreases in $r_{1}$.
Comparison of Schemes and Design Implications.
Having analyzed the effects of revert protection costs on searchers and the sequencer in both schemes, we now compare the two and draw implications for sequencer design.
Theorem 5.3
For $c$ sufficiently small, expected revenue in Scheme 2 is greater than the expected profit in Scheme 1 under the profit-optimal choice of $r_{1}$.
This result states that if the computational cost of implementing revert protection is sufficiently small, then having searchers pay the cost results in a greater expected profit to the sequencer than having the sequencer subsidize the costs itself. Besides yielding greater profits under small computational costs, the second scheme is also preferred in the sense that whereas the optimal $r_{1}$ needs to be adjusted based on auction parameters in the first scheme and is not necessarily zero, full revert protection is still optimal under the second scheme, with users paying a fixed cost rather than a penalty proportional to their bid, thus encouraging maximum participation in the PGA auction.
@@ Appendix 0.A Proofs
0.A.6 Proof of Theorem 5.1
The derivative of expected profit with respect to $r_{1}$ is
$\frac{\partial\mathbb{E}[\textsf{Profit}]}{\partial r_{1}}=-N\cdot\frac{\partial p^{*}}{\partial r_{1}}\left((p^{*})^{N-1}V-c\right)$
Since $N\cdot\frac{\partial p^{*}}{\partial r_{1}}>0$, dividing and equating to zero yields the FOC of
$(p^{*})^{N-1}V=c\implies r_{1}^{*}=\min\left\{\frac{c(V-g)}{(V-c)g},1\right\}\implies p^{*}=\left(\frac{c}{V}\right)^{\frac{1}{N-1}}.$
Since $\frac{\partial\mathbb{E}[\textsf{Profit}]}{\partial r_{1}}$ at $r_{1}=0$ is positive and the FOC has at most one solution over $r_{1}\in[0,1]$, it follows that $r_{1}^{*}$ is indeed a maximizer of expected profits.
0.A.7 Proof of Theorem 5.2
Similar to the proofs of Theorem 3.1, Theorem 3.2, and Theorem 3.3.
0.A.8 Proof of Theorem 5.3
Denote $\Pi_{1}$ and $\Pi_{2}$ as the expected profit in Scheme 1 under the optimal $r_{1}$ and expected revenue in Scheme 2 under $r_{1}=0$, respectively. It follows that
$\Pi_{1}=\left(1-\left(\frac{c}{V}\right)^{\frac{N}{N-1}}\right)V-c\left(\left(\frac{c}{V}\right)^{\frac{1}{N-1}}\right)N,\ \Pi_{2}=\left(1-\left(\frac{c}{V-g}\right)^{\frac{1}{N-1}}\right)V.$
Taking derivatives with respect to $c$ yields
$\frac{\partial\Pi_{1}}{\partial c}=-N\left(1-\left(\frac{c}{V}\right)^{\frac{1}{N-1}}\right),\ \frac{\partial^{2}\Pi_{1}}{\partial c^{2}}=\frac{1}{c(N-1)}\left(\frac{c}{V}\right)^{\frac{1}{N-1}},$
$\frac{\partial\Pi_{2}}{\partial c}=-\frac{VN}{(V-g)(N-1)}\left(\frac{c}{V-g}\right)^{\frac{1}{N-1}},\ \frac{\partial^{2}\Pi_{2}}{\partial c^{2}}=-\frac{1}{c(N-1)}\left(\frac{c}{V-g}\right)^{\frac{1}{N-1}}.$
Note that $\Pi_{1}=\Pi_{2}$ when $c=0$. Since $\Pi_{1}$ is decreasing and convex in $c$ while $\Pi_{2}$ is decreasing and concave in $c$, there exists at most one other $c$ at which $\Pi_{1}$ and $\Pi_{2}$ intersect. Since $\frac{\partial\Pi_{1}}{\partial c}|_{c=0}=-N$ and $\frac{\partial\Pi_{2}}{\partial c}|_{c=0}=0$, we have $\Pi_{2}\geq\Pi_{1}$ initially.